1. If a commercial bank holds 20% as required reserves, the initial deposit is ksh 10,000, what was the money created by the bank from this initial deposit?
1
Expert's answer
2022-03-07T18:08:56-0500

what was the money created by the bank from this initial deposit?

The reserve ratio is 20%. As a result, the money multiplier is "\frac{1}{20\%} = \frac{1}{20} = 5"


The money created is given by:

Created money = "\frac{Initial\ deposit }{Reserves\ ration }"


Created money = "\frac{10000 }{20\%}"


Created money = 50000

Thus, the money created is Ksh. 50000